OpenStudy (anonymous):

Easy Question :) Help though? (-1/2)x-(-3)=8

Nnesha (nnesha):

first multiply -1 by =3

OpenStudy (anonymous):

(-x/2) x -(-3) = 8 (-x/2)x (+3) = 8 [minus by minus = +] (-x/2)x 3 = 8 [numerator by numerator] -3x/2 = 8 -3x = 8 x 2 [ multiply both sides by 2] -3x = 16 3x = -16 [ multiply both sides by -1] x = -16/3 or -5 and 1/3.

Nnesha (nnesha):

first of multiply -1 by -3 then you have left +3 3 is adding on left side not multiplying by -1/2 x